# Holistic Unlearning Benchmark: A Multi-Faceted Evaluation for Text-to-Image Diffusion Model Unlearning

### v2

@@ 3.2.5 Attack robustness
The attack robustness perspective evaluates whether unlearning methods are robust against optimization-based attacks. We employ Ring-a-Bell [42] to generate 1,000 prompts per concept. Specifically, Ring-a-Bell optimizes a randomly initialized prompt using a CLIP text encoder to align closely with the target concept word in the CLIP space. We report the proportion of the target concepts in images generated with the optimized prompts.
@@ 4 Benchmark Results
We conduct experiments on 15 tasks identified across six different perspectives for seven different unlearning methods: $\mathtt{SLD}$ [38], $\mathtt{AC}$ [20], $\mathtt{ESD}$ [9], $\mathtt{UCE}$ [10], $\mathtt{SA}$ [13], $\mathtt{RECELER}$ [16], and $\mathtt{MACE}$ [24]. We use Stable Diffusion v1.5 [34] as the original model of our evaluation. The detailed explanations and configurations of the baselines are presented in Appendix D. Overall evaluation results are presented in Tab. 3. Detailed results are explained below.
@@
It is worth noting that the original model may appear to perform lower than the unlearning methods for selective alignment. We hypothesize that this occurs since the original model inherently incorporates the target concept into its generated images, adding an extra compositional element that makes it difficult for the model to generate the remaining concepts [23]. For reference, we retain the selective alignment performance of the original model.
@@ 5 Analysis and Discussion
Comparison between I2P dataset and our dataset. Many unlearning methods evaluate the unlearning performance of NSFW concepts through the I2P dataset [38], known as a collection of inappropriate prompts. However, as reported in the dataset [38], only 2.8% of images have a NudeNet [2] probability over 50%, and only 37.2% are classified as inappropriate by Q16 [37], raising concern about I2P prompts as a benchmark for $\mathtt{NSFW}$ generation. In Tab. 7, we compare the target proportions in generated images with I2P prompts and our prompts, respectively. The results suggest that our newly curated prompts can generate more $\mathtt{NSFW}$-relevant images than I2P.
@@ A.2 Prompt Generation
For the prompt generation framework, we utilize Meta-Llama-3.1-8B-Instruct [11] for the $\mathtt{Celebrity}$, $\mathtt{Style}$, and $\mathtt{IP}$. For $\mathtt{NSFW}$, we use Meta-Llama-3.1-8B-Instruct-abliterated [21].
@@ B.1 Prompt Instruction for the Framework
1) Reference Images (A $<$image$>$, B $<$image$>$, C $<$image$>$):
@@
2) Query Image (D $<$image$>$):
@@ B.2 Verification of the Framework
To demonstrate the effectiveness of our proposed VLM-based concept detection framework, we conduct extensive evaluations using two datasets: the Disney character dataset [36] for $\mathtt{IP}$, and AI-ArtBench [41] for $\mathtt{Style}$. Specifically, the Disney character dataset contains images from five Disney characters: Mickey Mouse, Donald Duck, Minions, Olaf, and Pooh, with approximately 90 test images per character. The AI-ArtBench dataset consists of images generated using a diffusion model, with 1,000 images per artistic style. In our evaluation, we select five artistic styles: Expressionism, Impressionism, Renaissance, Surrealism, and Ukiyo-e.
We evaluate our concept detection framework using two VLMs, InternVL2.5-8B [4] and Qwen2.5-VL-7B [1], as backbone models. For each VLM backbone, we measure the true positive rate (TPR) and the false positive rate (FPR) as metrics for detection accuracy. For $\mathtt{Style}$, InternVL2.5-8B achieves an average TPR of 82.5% and an average FPR of 4.7%. With Qwen2.5-VL-7B, TPR and FPR are 85.1% and 0.5%, respectively. For $\mathtt{IP}$, InternVL2.5-8B achieves a TPR of 83.2% and a FPR of 1.5%, while Qwen2.5-VL-7B achieves a TPR of 85.1% and a FPR of 0.5%.
@@
Table 14: Experimental results on VLM concept detection. We evaluate our VLM concept detection method on two datasets: one with $\mathtt{Style}$ images and another with $\mathtt{IP}$ images. We measure the true positive rate (TPR) and the false positive rate (FPR) for five representative concepts in each dataset. We utilize InternVL2.5-8B [4] and Qwen2.5-VL-7B [1] as the backbone models for our VLM-based concept detection framework.
@@ C.2 Pinpoint-ness
Appendix D Baselines
In this section, we describe the baseline unlearning methods and training settings used in our evaluation. For all experiments, we use Stable Diffusion v1.5 [34] as the original text-to-image diffusion model. For $\mathtt{Celebrity}$, $\mathtt{Style}$, and $\mathtt{IP}$, we individually train separate unlearning models, each specialized to remove a single specific concept. In contrast, for $\mathtt{NSFW}$, we train a single unified model to simultaneously unlearn all three categories: Nudity, Disturbing, and Violent.
@@ D.1 Safe Latent Diffusion (SLD)
$\mathtt{SLD}$ [38] mitigates the generation of images containing a target concept by incorporating a negative prompt. Specifically, during classifier-free guidance, the diffusion model utilizes outputs conditioned on this negative prompt to guide the image generation process away from undesired content. $\mathtt{SLD}$ is categorized into four variants, SLD-Weak, SLD-Medium, SLD-Strong, and SLD-Max, depending on the hyperparameter settings controlling the strength of unlearning. For our experiments, we adopt the SLD-Medium variant. We use the target concepts directly as negative prompts for $\mathtt{Celebrity}$, $\mathtt{Style}$, and $\mathtt{IP}$. For the $\mathtt{NSFW}$ benchmark, we follow the original implementation by using the same predefined set of negative prompts.
@@ D.2 Ablating Concept (AC)
$\mathtt{AC}$ [20] employs an alternative concept $c^{*}$ to prevent the generation of a specific target concept $c$. The objective is defined as follows:
@@
For $\mathtt{Style}$ and $\mathtt{NSFW}$, we adopt the experimental settings from the original implementation. For $\mathtt{IP}$, we select “animated character” as the alternative concept. For $\mathtt{Celebrity}$, we use “middle aged man (woman)” as the alternative concept, consistent with the setting in $\mathtt{SA}$, and additionally increase the number of training steps to 400 and the learning rate to 4e-6.
@@ D.3 Selective Amnesia (SA)
$\mathtt{SA}$ [13] leverages techniques from continual learning, including Elastic Weight Consolidation (EWC) [18] and generative replay (GR) [40]:
@@
For $\mathtt{Style}$, $\mathtt{Celebrity}$, and $\mathtt{IP}$, we use “painting”, “middle aged man (woman)”, and “animated character” as the alternative concepts, respectively. For $\mathtt{NSFW}$, we employ “people” as the alternative concept, as in $\mathtt{AC}$. All other hyperparameters remain unchanged from the original implementation.
@@ D.4 Erased Stable Diffusion (ESD)
$\mathtt{ESD}$ [9] fine-tunes the diffusion model using the following objective:
@@
where $\theta$ denotes the trainable parameters of the diffusion model, $\theta^{*}$ represents the fixed original diffusion model, and $c$ represents the target concept. Intuitively, this modified score function shifts the learned data distribution away from the target concept $\boldsymbol{c}$, thereby reducing the likelihood of generating images containing the undesired concept.
@@ D.5 Unified Concept Editing (UCE)
$\mathtt{UCE}$ [10] edit weights of cross-attention layers for its unlearning:
@@
The destination embedding for the object unlearning is equal to a null embedding (i.e., “”). For our evaluation, we employ the same training settings as in the original implementation.
@@ D.6 Reliable Concept Erasing via Lightweight Erasers (RECELER)
$\mathtt{RECELER}$ [16] trains an adapter-based eraser $E$ by employing the same objective as $\mathtt{ESD}$. Additionally, $\mathtt{RECELER}$ incorporates a masking-based regularization loss, which encourages the eraser to selectively remove only the specified target concept. For our experiments, we utilize the original implementation of $\mathtt{RECELER}$ to train models. All training configurations in our evaluation exactly follow the settings from the original implementation.
@@ D.7 Mass Concept Erasure (MACE)
$\mathtt{MACE}$ [24] computes an attention map for a given input image by utilizing the cross-attention layers of the diffusion model, and leverages this attention map to facilitate the unlearning of the target concept. Additionally, $\mathtt{MACE}$ employs a Low-Rank Adaptation (LoRA) module to efficiently fine-tune the diffusion model. In our experiments, we follow the original training configuration from the $\mathtt{MACE}$ implementation, except for $\mathtt{NSFW}$. Specifically, for $\mathtt{NSFW}$, we train the diffusion model using two distinct sets of concept words: (1) the original word set from the $\mathtt{MACE}$ implementation, and (2) the concept word list previously employed by $\mathtt{AC}$ for $\mathtt{NSFW}$ unlearning.
@@ E.1 Target Proportion
$\mathtt{Celebrity}$ Angelina Jolie Ariana Grande Brad Pitt David Beckham Elon Musk Emma Watson Lady Gaga Leonardo DiCaprio Taylor Swift Tom Cruise Average
@@
$\mathtt{Style}$ Andy Warhol Auguste Renoir Claude Monet Frida Kahlo Paul Cézanne Pablo Picasso Piet Mondrian Roy Lichtenstein Van Gogh Édouard Manet Average
@@
$\mathtt{IP}$ Buzz Lightyear Homer Simpson Luigi Mario Mickey Mouse Pikachu Snoopy Sonic SpongeBob Stitch Average
@@ E.2 General Image Quality
$\mathtt{Celebrity}$ Angelina Jolie Ariana Grande Brad Pitt David Beckham Elon Musk Emma Watson Lady Gaga Leonardo DiCaprio Taylor Swift Tom Cruise Average
@@
$\mathtt{Style}$ Andy Warhol Auguste Renoir Claude Monet Frida Kahlo Paul Cézanne Pablo Picasso Piet Mondrian Roy Lichtenstein Van Gogh Édouard Manet Average
@@
$\mathtt{IP}$ Buzz Lightyear Homer Simpson Luigi Mario Mickey Mouse Pikachu Snoopy Sonic SpongeBob Stitch Average
@@
$\mathtt{Celebrity}$ Angelina Jolie Ariana Grande Brad Pitt David Beckham Elon Musk Emma Watson Lady Gaga Leonardo DiCaprio Taylor Swift Tom Cruise Average
@@
$\mathtt{Style}$ Andy Warhol Auguste Renoir Claude Monet Frida Kahlo Paul Cézanne Pablo Picasso Piet Mondrian Roy Lichtenstein Van Gogh Édouard Manet Average
@@
$\mathtt{IP}$ Buzz Lightyear Homer Simpson Luigi Mario Mickey Mouse Pikachu Snoopy Sonic SpongeBob Stitch Average

### v3

@@ 3.2.5 Attack robustness
The attack robustness perspective evaluates whether unlearning methods are robust against optimization-based attacks. We employ Ring-a-Bell [ringabell] to generate 1,000 prompts per concept. Specifically, Ring-A-Bell optimizes a randomly initialized prompt using a CLIP text encoder to align closely with the target concept word in the CLIP space. We report the proportion of the target concepts in images generated with the optimized prompts. Additionally, to further assess robustness, we include evaluations using two more attacks: UnlearnDiffAtk (UDA) [zhang2024generate] and Unlearning or Concealment (UoC) [sharma2024unlearning]. Detailed descriptions and results for the attacks are provided in the Sec. E.8.
@@ 4 Benchmark Results
We conduct experiments on 15 tasks identified across six different perspectives for seven different unlearning methods: $\mathtt{SLD}$ [schramowski2023safe], $\mathtt{AC}$ [kumari2023ablating], $\mathtt{ESD}$ [gandikota2023erasing], $\mathtt{UCE}$ [gandikota2024unified], $\mathtt{SA}$ [heng2024selective], $\mathtt{RECELER}$ [huang2023receler], and $\mathtt{MACE}$ [lu2024mace]. We use Stable Diffusion v1.5 [rombach2022high] as the original model of our evaluation. Detailed explanations and training configurations of the baselines are presented in Appendix D. The evaluation results are presented in Tab. 3. Detailed results are explained below.
@@
It is worth noting that the original model may appear to perform lower than the unlearning methods for selective alignment. We hypothesize that this occurs since the original model inherently incorporates the target concept into its generated images, adding an extra compositional element that makes it difficult for the model to generate the remaining concepts [liu2022compositional]. For reference, we retain the selective alignment performance of the original model.
@@ 5 Analysis and Discussion
Comparison between I2P dataset and our dataset. Many unlearning methods evaluate the unlearning performance of NSFW concepts through the I2P dataset [schramowski2023safe], known as a collection of inappropriate prompts. However, as reported in the dataset [schramowski2023safe], only 2.8% of images have a NudeNet [nudenet] probability over 50%, and only 37.2% are classified as inappropriate by Q16 [schramowski2022can], raising concern about I2P prompts as a benchmark for $\mathtt{NSFW}$ generation. In Tab. 7, we compare the target proportions in generated images with I2P prompts and our prompts, respectively. The results suggest that our newly curated prompts can generate more $\mathtt{NSFW}$-relevant images than I2P.
@@ 6 Conclusion
Acknowledgements
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) and the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (RS-2019-II191906, Artificial Intelligence Graduate School Program (POSTECH); RS-2024-00457882, National AI Research Lab Project; RS-2024-00509258 and RS-2024-00469482, Global AI Frontier Lab; RS-2025-00560062; RS-2023-00217286). This research was also supported by Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (RS-2024-00406787) and the Hyundai Motor Chung Mong-Koo Foundation.
@@ A.2 Prompt Generation
For the prompt generation framework, we utilize Meta-Llama-3.1-8B-Instruct [grattafiori2024llama] for the $\mathtt{Celebrity}$, $\mathtt{Style}$, and $\mathtt{IP}$. For $\mathtt{NSFW}$, we use Meta-Llama-3.1-8B-Instruct-abliterated [llama_3_1_8b_abliterated].
@@
A.3 Statistical Analysis.
To assess statistical stability, we repeat each experiment five times with different seeds on the $\mathtt{NSFW}$ category. As shown in Tab. 13, the standard deviation remains below 3% across all methods and metrics, indicating consistent results.
Target proportion Pinpoint-ness Multilingual robustness Attack robustness
$\mathtt{Original}$ $0.647_{\pm{0.004}}$ $0.589_{\pm{0.015}}$ $0.324_{\pm{0.010}}$ $0.801_{\pm{0.003}}$
$\mathtt{SLD}$ $0.340_{\pm{0.003}}$ $527_{\pm{0.016}} 0.0$$079_{\pm{0.001}}$ $0.496_{\pm{0.008}}$
$\mathtt{AC}$ $0.439_{\pm{0.002}}$ $0.467_{\pm{0.012}}$ $0.177_{\pm{0.003}}$ $0.542_{\pm{0.026}}$
$\mathtt{ESD}$ $0.344_{\pm{0.001}}$ $0.127_{\pm{0.012}}$ $0.159_{\pm{0.003}}$ $0.479_{\pm{0.007}}$
$\mathtt{UCE}$ $0.604_{\pm{0.001}}$ $0.550_{\pm{0.015}}$ $0.252_{\pm{0.003}}$ $0.776_{\pm{0.006}}$
$\mathtt{SA}$ $0.330_{\pm{0.001}}$ $0.102_{\pm{0.009}}$ $0.125_{\pm{0.002}}$ $0.450_{\pm{0.008}}$
$\mathtt{RECELER}$ $0.271_{\pm{0.002}}$ $306_{\pm{0.013}} 0.0$$082_{\pm{0.001}}$ $0.365_{\pm{0.014}}$
$\mathtt{MACE}$ $0.342_{\pm{0.002}}$ $0.126_{\pm{0.016}}$ $0.348_{\pm{0.005}}$ $0.358_{\pm{0.005}}$
Table 13: Evaluation results on the $\mathtt{NSFW}$ category along with the corresponding standard deviations.
@@ B.1 Prompt Instruction for the Framework
1) Reference Images (A <<image>>, B <<image>>, C <<image>>):
@@
2) Query Image (D <<image>>):
@@ B.2 Verification of the Framework
To demonstrate the effectiveness of our proposed VLM-based concept detection framework, we conduct extensive evaluations using two datasets: the Disney character dataset [disney_characters_dataset] for $\mathtt{IP}$, and AI-ArtBench [ai_artbench] for $\mathtt{Style}$. Specifically, the Disney character dataset contains images from five Disney characters: Mickey Mouse, Donald Duck, Minions, Olaf, and Pooh, with approximately 90 test images per character. The AI-ArtBench dataset consists of images generated using a diffusion model, with 1,000 images per artistic style. In our evaluation, we select five artistic styles: Expressionism, Impressionism, Renaissance, Surrealism, and Ukiyo-e.
We evaluate our concept detection framework using two VLMs, InternVL2.5-8B [chen2024expanding] and Qwen2.5-VL-7B [bai2025qwen2], as backbone models. For each VLM backbone, we measure the true positive rate (TPR) and the false positive rate (FPR) as metrics for detection accuracy. For $\mathtt{Style}$, InternVL2.5-8B achieves an average TPR of 82.5% and an average FPR of 4.7%. With Qwen2.5-VL-7B, TPR and FPR are 85.1% and 0.5%, respectively. For $\mathtt{IP}$, InternVL2.5-8B achieves a TPR of 83.2% and a FPR of 1.5%, while Qwen2.5-VL-7B achieves a TPR of 85.1% and a FPR of 0.5%.
@@
Table 15: Experimental results on VLM concept detection. We evaluate our VLM concept detection method on two datasets: one with $\mathtt{Style}$ images and another with $\mathtt{IP}$ images. We measure the true positive rate (TPR) and the false positive rate (FPR) for five representative concepts in each dataset. We utilize InternVL2.5-8B [chen2024expanding] and Qwen2.5-VL-7B [bai2025qwen2] as the backbone models for our VLM-based concept detection framework.
@@ C.2 Pinpoint-ness
Appendix D Baselines and Training Details
For all experiments, we use Stable Diffusion v1.5 [rombach2022high] as the original text-to-image diffusion model. For $\mathtt{Celebrity}$, $\mathtt{Style}$, and $\mathtt{IP}$, we individually train separate unlearning models, each specialized to remove a single specific concept. For $\mathtt{NSFW}$, we train a model to simultaneously unlearn all three categories: Nudity, Disturbing, and Violent.
@@ D.1 Safe Latent Diffusion (SLD)
$\mathtt{SLD}$ [schramowski2023safe] mitigates the generation of images containing a target concept by incorporating a negative prompt. Specifically, during classifier-free guidance, the diffusion model utilizes outputs conditioned on this negative prompt to guide the image generation process away from undesired content. $\mathtt{SLD}$ is categorized into four variants, SLD-Weak, SLD-Medium, SLD-Strong, and SLD-Max, depending on the hyperparameter settings controlling the strength of unlearning. We use SLD-Medium for all experiments. We use the target concepts directly as negative prompts for $\mathtt{Celebrity}$, $\mathtt{Style}$, and $\mathtt{IP}$. For $\mathtt{NSFW}$, we follow the original implementation by using the same predefined set of negative prompts.
@@ D.2 Ablating Concept (AC)
$\mathtt{AC}$ [kumari2023ablating] employs an alternative concept $c^{*}$ to prevent the generation of a specific target concept $c$. The objective is defined as follows:
@@
We adopt the experimental settings from the original implementation. For $\mathtt{IP}$, $\mathtt{Style}$, and $\mathtt{Celebrity}$, we use “animated character”, “painting”, and “middle aged man (woman)” as the alternative concept, respectively. For $\mathtt{IP}$, we use the prompts used in the original implementation. For the remaining hyperparameters, such as the number of training steps and learning rate, we use the hyperparameters used in the original implementation, except for $\mathtt{Celebrity}$. For $\mathtt{Celebrity}$, we increase the number of training steps to 400 and the learning rate to 4e-6, resulting in improved performance.
@@ D.3 Selective Amnesia (SA)
$\mathtt{SA}$ [heng2024selective] leverages techniques from continual learning, including Elastic Weight Consolidation (EWC) [kirkpatrick2017overcoming] and generative replay (GR) [shin2017continual]:
@@
For $\mathtt{Style}$, $\mathtt{Celebrity}$, and $\mathtt{IP}$, we use “painting”, “middle aged man (woman)”, and “animated character” as the alternative concepts, respectively, as in $\mathtt{AC}$. For $\mathtt{NSFW}$, we employ “people” as the alternative concept. All other hyperparameters remain unchanged from the original implementation. Specifically, we use 200 epochs for $\mathtt{Style}$, $\mathtt{Celebrity}$, and $\mathtt{IP}$, and 500 epochs for $\mathtt{NSFW}$. We also set the learning rate to 1e-5.
@@ D.4 Erased Stable Diffusion (ESD)
$\mathtt{ESD}$ [gandikota2023erasing] fine-tunes the diffusion model using the following objective:
@@
where $\theta$ denotes the trainable parameters of the diffusion model, $\theta^{*}$ represents the fixed original diffusion model, and $c$ represents the target concept. Intuitively, this modified score function shifts the learned data distribution away from the target concept $\boldsymbol{c}$, thereby reducing the likelihood of generating images containing the undesired concept. $\mathtt{ESD}$ can be categorized into ESD-x and ESD-u, depending on which parameters are fine-tuned. ESD-x fine-tunes only the cross-attention parameters in the U-Net, while ESD-u updates only the unconditional parameters of the U-Net.
For training, we use the same hyperparameters used in the original implementation. Following the original paper, we apply ESD-x to $\mathtt{Style}$, $\mathtt{Celebrity}$, and $\mathtt{IP}$. For $\mathtt{NSFW}$, we use ESD-u instead. We use 200 training steps and set the learning rate to 5e-5 for the training.
@@ D.5 Unified Concept Editing (UCE)
$\mathtt{UCE}$ [gandikota2024unified] edit weights of cross-attention layers for its unlearning:
@@
The destination embedding for the object unlearning is equal to a null embedding (i.e., “”). For training, we employ the same training settings as in the original implementation. We use the target concepts directly as negative prompts for $\mathtt{Celebrity}$, $\mathtt{Style}$, and $\mathtt{IP}$. For $\mathtt{NSFW}$, we follow the original implementation by using the same predefined set of negative prompts.
@@ D.6 Reliable Concept Erasing via Lightweight Erasers (RECELER)
$\mathtt{RECELER}$ [huang2023receler] trains an adapter-based eraser $E$ by employing the same objective as $\mathtt{ESD}$. Additionally, $\mathtt{RECELER}$ incorporates a masking-based regularization loss, which encourages the eraser to selectively remove only the specified target concept. For our experiments, we utilize the original implementation of $\mathtt{RECELER}$ to train models. All training configurations in our evaluation exactly follow the settings from the original implementation. We set the training steps to 500 and learning rate to 3e-4, respectively.
@@ D.7 Mass Concept Erasure (MACE)
$\mathtt{MACE}$ [lu2024mace] computes an attention map for a given input image by utilizing the cross-attention layers of the diffusion model, and leverages this attention map to facilitate the unlearning of the target concept. $\mathtt{MACE}$ leverages the Grounded SAM to generate segmentation masks indicating the location of the target concept in the given training images, which are then used for unlearning. $\mathtt{MACE}$ additionally employs a Low-Rank Adaptation (LoRA) module to efficiently fine-tune the diffusion model. In our experiments, we generally follow the original training configuration from the $\mathtt{MACE}$ implementation. Specifically, we set the learning rate to 1e-4 and max training step to 50 for $\mathtt{IP}$, $\mathtt{Style}$, and $\mathtt{Celebrity}$. For $\mathtt{NSFW}$we set the learning rate to 1e-5 and max training step to 120. For $\mathtt{NSFW}$. $\mathtt{IP}$, $\mathtt{Style}$, $\mathtt{Celebrity}$, and $\mathtt{NSFW}$, we use “animated character”, “art”, “man (woman)”, and “person” as the alternative concept, respectively.
@@ E.1 Target Proportion
$\mathtt{Celebrity}$
Angelina Jolie Ariana Grande Brad Pitt David Beckham Elon Musk Emma Watson Lady Gaga Leonardo DiCaprio Taylor Swift Tom Cruise Average
@@
$\mathtt{Style}$
Andy Warhol Auguste Renoir Claude Monet Frida Kahlo Paul Cézanne Pablo Picasso Piet Mondrian Roy Lichtenstein Van Gogh Édouard Manet Average
@@
$\mathtt{IP}$
Buzz Lightyear Homer Simpson Luigi Mario Mickey Mouse Pikachu Snoopy Sonic SpongeBob Stitch Average
@@ E.2 General Image Quality
$\mathtt{Celebrity}$
Angelina Jolie Ariana Grande Brad Pitt David Beckham Elon Musk Emma Watson Lady Gaga Leonardo DiCaprio Taylor Swift Tom Cruise Average
@@
$\mathtt{Style}$
Andy Warhol Auguste Renoir Claude Monet Frida Kahlo Paul Cézanne Pablo Picasso Piet Mondrian Roy Lichtenstein Van Gogh Édouard Manet Average
@@
$\mathtt{IP}$
Buzz Lightyear Homer Simpson Luigi Mario Mickey Mouse Pikachu Snoopy Sonic SpongeBob Stitch Average
@@
$\mathtt{Celebrity}$
Angelina Jolie Ariana Grande Brad Pitt David Beckham Elon Musk Emma Watson Lady Gaga Leonardo DiCaprio Taylor Swift Tom Cruise Average
@@
$\mathtt{Style}$
Andy Warhol Auguste Renoir Claude Monet Frida Kahlo Paul Cézanne Pablo Picasso Piet Mondrian Roy Lichtenstein Van Gogh Édouard Manet Average
@@
$\mathtt{IP}$
Buzz Lightyear Homer Simpson Luigi Mario Mickey Mouse Pikachu Snoopy Sonic SpongeBob Stitch Average